OpenStudy (anonymous):

A hockey player makes a slap shot, exerting a force of 42.5 N on the hockey puck for 0.168 s. What impulse is given to the puck?

OpenStudy (anonymous):

multiply those two together and you should get your answer since F(t)=p=mv2-mv1 and since you need p just do p=F(t)

OpenStudy (anonymous):

I=F*t=42.5*0.168=7.14

OpenStudy (anonymous):

also it would be in N*s

OpenStudy (anonymous):

so 7.14 N*s

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you